OpenStudy (anonymous):

Find the x-intercepts (if any) for the graph of the quadratic function. 6x2 + 12x + 5 = 0 Give your answers in exact form. Show your work.

hartnn (hartnn):

to get x intercepts plug in y=0

hartnn (hartnn):

ok, you don't have a y there just solve that quadratic :P

OpenStudy (anonymous):

i did i check it

hartnn (hartnn):

sure

OpenStudy (anonymous):

she ask me exact form i dontknow why its not coming maybe my calculator error plzz re-d and check

hartnn (hartnn):

it has x intercepts

OpenStudy (anonymous):

check my attacthment

hartnn (hartnn):

\((-12\times 5)\) ^^ incorrect

hartnn (hartnn):

shouldn't it just be 12*5 ?

OpenStudy (anonymous):

@master50777 u here?

OpenStudy (anonymous):

oh yes @hartnn next error?

hartnn (hartnn):

no other error :) answer will change

OpenStudy (anonymous):

but formal had negative sign already

hartnn (hartnn):

\(b^2-4ac \\ 12^2 -4(12)(5)\)

OpenStudy (anonymous):

144-56

OpenStudy (anonymous):

Sqrt 88

hartnn (hartnn):

sorry \(12^2-4\times 6\times 5\)

hartnn (hartnn):

144-120 = 24 \(\sqrt {24}\)

OpenStudy (anonymous):

final answer 0.41

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

ok, \(\dfrac{-12 \pm \sqrt{24}}{12} = \dfrac{-3\pm \sqrt 6}{3}\)

hartnn (hartnn):

yeah, correct :)

OpenStudy (anonymous):

thnaks :)

OpenStudy (anonymous):

ill be doing my other test which is complete but has some error and i need to review could u help me in those Q's

hartnn (hartnn):

sure :) just ask each question in a new post :)

OpenStudy (anonymous):

ok :)